# Membrane Paradigm Approach to the Johannsen-Psaltis Black Hole

### v1

@@ Abstract
The Johannsen-Psaltis black hole is a phenomenologically viable metric obtained by judiciously deforming the Kerr black hole such that the metric is asymptotically flat and causal; and is consistent with the weak field, post-Newtonian tests of gravity; but has additional hairs besides its mass and angular momentum. As it deviates from the Kerr black hole in the strong coupling regime, it is a useful metric to test the Kerr Hypothesis that states that all astrophysical black holes are of the Kerr type. Here we give a membrane description of this black hole that effectively amounts to replacing the observable part of the black hole with a fluid with thermal properties. A time-like membrane, a stretched horizon local in time is assumed to exist. This membrane is expected to approximate the null event horizon which is highly non-local in time. We derive the energy-momentum tensor of the fluid and all the transport coefficients using the action formulation to the membrane as advocated by Parikh and Wilczek. In the fluid description, one observes that the finiteness of the transport coefficients constrains the additional hairs of the Johannsen-Psaltis black hole. Analytically continuing the pressure of the fluid to all values of the radial coordinate $r$, one obtains interesting Van der Waals-type behavior of the pressure of the fluid which diverges at the radius of the outer ergosphere, lending support, from the Membrane Paradigm’s perspective to the claim that relativistic astrophysical jets are produced by the ergoregion of the black hole.
@@ I Introduction
Most of the deformations of the Kerr metric yield pathological results such as the existence of closed time-like curves, or naked singularities. Moreover, because of the powerful uniqueness theorems of the Kerr black hole [8] within General Relativity, it is just in vain to modify the Kerr metric as a solution to Einstein’s equations: one must either consider alternative theories of gravity that allow viable deformations of the Kerr metric in the strong field regime or, perhaps better yet, without referring to any particular field equations, consider phenomenological metrics that judiciously deform metric. In the literature, one can find many works devoted to the first type of deformations, that is modified Kerr solutions (albeit almost always approximate solutions) to modified theories of Kerr gravity, and not so many phenomenological metrics except the ones constructed by Johannsen and Psaltis [10, 9]. Here we shall study these metrics in the context of the membrane paradigm approach to black holes which is an effective description of the black hole in terms of a fluid with various thermal properties. This is an exotic fluid (which may not be easily drinkable as it has a negative bulk viscosity) that mimics the true Event Horizon of the black hole. Our motivation to study the modified Kerr metric in the membrane paradigm was to understand if there are some constraints on the deformations parameters, and indeed there are as we shall see. In some sense, the theory is still a great "experimental tool" and a viable membrane paradigm that was quite useful in understanding various aspects of black holes in General Relativity, helping us constrain the deformation parameter, indeed a new hair, in the parametrically deformed Kerr metric.
The layout of the paper is as follows: In section II we give a detailed account of the membrane construction using the action formalism of Parikh-Wilczek [11]; Section III is the bulk of our paper: we construct the membranes for JP metric after considering he membrane for its static limit. In the Appendix, we discuss an earlier version of the JP metric and its membrane description. The computations presented here can be applied verbatim to any given stationary metric that forms an asymptotically flat deformation of the Kerr black hole.
@@ II.1 Basics of the Membrane Paradigm
The event horizon of a static or stationary black hole is a null hypersurface of codimension one with a degenerate metric. It is highly non-local in time and hence not observable to transient observers like us and is not a tangible concept for astrophysical purposes. To remedy this and to approximate the black hole as a local object in time, an effective description that makes use of time-like surfaces, not null surfaces is needed. For this purpose, a "stretched horizon" or "fluid bubble", that is a time-like surface [12] with electrical conductivity having a resistance of 377 $\Omega$, shear &\&& bulk viscosities [13] arbitrarily close to the event horizon was introduced by Damour [13]. This pragmatic point of view, that is endowing a black hole with a time-like facade has proven to be very useful. A careful calculation on the horizon led to Ohm’s law [14], Joule’s law, and the non-relativistic Navier-Stokes equation [15]. Early works of Damour, followed by Thorne and Price [16] showed a way to mimic the null horizon in a well-defined approximate way since the membrane is not null by nature, it has a non-degenerate Lorentzian metric which allows computations local in time. This approach to black hole dynamics was coined as the Membrane Paradigm of Black Holes [16]. More recently [11], a proper action formulation of the membrane was given that makes the computations rather straightforward, and hence we shall use this. See [17] for a recent review of these ideas.
@@ II.2 Quantities Describing the Fluid Membrane
Let $\mathcal{H}$ be the event horizon of a stationary black hole, then there exists, by definition, a null geodesic vector field $\ell$ that generates $\mathcal{H}$. One can define a time-like stretched horizon $\mathcal{H}_{s}$ arbitrarily close to this event horizon. In the ADM decomposition of the metric [18], let $N$ be the lapse function which can be chosen such that in the limit $N\rightarrow 0$, the stretched horizon goes to the true horizon, that is $\mathcal{H}_{s}|_{N\rightarrow 0}=\mathcal{H}$ [15]. This limit in the geometry cannot be smooth and hence $N$ will be used as a regulator in various geometric objects as we shall see below.
Let $(\mathcal{M},g)$ be the $(3+1)$-dimensional total black hole spacetime and $(\mathcal{H}_{s},h)$ be a $(2+1)$-dimensional submanifold of $(\mathcal{M},g)$ with the pull-back metric $h=\phi^{*}g$, which can also be considered as the projector $h^{\mu}\,_{\nu}:T_{p}(\mathcal{M})\longrightarrow T_{p}(\mathcal{H}_{s})$. A space-like cross-section of $\mathcal{H}_{s}$ is also a submanifold of dimension 2, which we denote as $(\Sigma,\gamma)$. We can do differential geometry adapted to these two submanifolds which boils down to a $(2+1+1)$-splitting of the full spacetime. Let $\nabla_{\mu}$ be the $g$-compatible covariant derivative and $D_{\mu}$ be the $h$-compatible covariant derivative while $\mathcal{D}_{\mu}$ be the $\gamma$-compatible covariant derivative [11]. Then, let $V^{\mu}\in T_{p}(\mathcal{M})$ and $n^{\mu}$ be a space-like unit vector normal to $\mathcal{H}_{s}$, then defining the extrinsic curvature of $\mathcal{H}_{s}$ as $K_{\sigma\mu}:={h}{{}^{\sigma}}_{\mu}\nabla_{\sigma}n^{\mu}$, one has the identity
@@
Let $u^{\mu}$ be the unit normal time-like vector to $\mathcal{H}_{s}$ chosen to satisfy $u^{\mu}n_{\mu}=0$. $u^{\mu}$ can be considered to be the four-velocity of a fiducial observer with proper time $\tau$. Here is the crucial part of the discussion: for the membrane paradigm to approximate the true horizon, a nullness constraint on the stretched horizon should be imposed; that means a change of character of $u^{\mu}$ and $n^{\mu}$ in the limits
@@
The relations we shall use for the Membrane Paradigm can be summarized as [11]:
@@ II.3 An Action Formalism for the Membrane Paradigm
Parikh and Wilczek [11] gave an action formulation of the membrane which starts by modifying the usual variational principle
@@
where the surface term refers to the black hole boundary which now is represented effectively by the membrane. The first and second parts of the action variations are assumed to be zero individually [15].
@@ II.3.1 The Gravitational Membrane
where we assumed $G_{N}=1$, $c=1$ and the second term is the Gibbons-Hawking boundary term. Using the Palatini’s identity, one has [19]:
@@
We choose the normal unit vector $n^{\mu}$ as outward-pointing. Applying the Leibniz rule to the integrand gives:
@@
Using the definition of the extrinsic curvature as given in the paragraph above (1), one finds the variation of the action as the Brown-York quasi-local stress tensor.
@@
Then, one has $t^{\text{stretched}}_{\mu\nu}=\frac{1}{8\pi}\left(Kh_{\mu\nu}-K_{\mu\nu}\right)\in T_{p}(\mathcal{H}_{s})\otimes T_{p}(\mathcal{H}_{s})$. On the stretched horizon $g^{\mu\nu}|_{\mathcal{H}_{s}}=h^{\mu\nu}$.To cancel the above non-zero boundary variation term, one must add the following:
@@
As it happens in electrodynamics where a surface charge induces a discontinuity in the field strength on the surface, $t^{\text{stretched}}_{\mu\nu}$ induces a discontinuity in the stretched horizon’s extrinsic curvature $K_{\mu\nu}$ [11]. This discontinuity creates a junction on the surface which can be identified as the Israel junction condition [20].
@@
Hence the gravitational membrane acts like a fluid obeying Damour-Navier-Stokes equations on the space-like cross-section of the stretched horizon [21]. If one can write ${K}{{}^{\mu}}_{\nu}$ in terms of the surface gravity $\kappa$ and extrinsic curvature ${k}{{}^{A}}_{B}$ of the space-like section of $\mathcal{H}_{s}$. To this end, one has $\nabla_{\ell}\ell=\kappa_{r}\ell$ where $\kappa_{r}$ is the normalized surface gravity at the horizon, which is related to the surface gravity $\kappa$ as $\kappa_{r}=N\kappa$ [11].
@@
where $\sigma_{AB}$ is the shear tensor. Then finally, the stretched horizon stress tensor becomes [15]:
@@ III.1 A membrane for the Schwarzschild-type geometry
where $f=1-\frac{2m(r)}{r}$. we can the metric as: [^1] This construction follows [22].
@@
of which the trace is $K=\frac{1}{2}\frac{\partial_{r}f}{\sqrt{f}}+2\frac{\sqrt{f}}{r}$. In the $(2+1+1)$-splitting of the spacetime, the extrinsic curvature on $\mathcal{H}_{s}$ can be identified by choosing the lapse function $N=\sqrt{f}$ as a renormalization factor:
@@
where $k_{\mu\nu}=\gamma_{\mu A}\gamma_{\nu B}k^{AB}$ is the extrinsic curvature of the 2D surface; and $\kappa$ is the surface gravity [22].
@@
equations (19), (20),(21) can be combined to find the stress tensor in terms of given parameters:
@@
with energy density $\rho$, pressure $P$, null geodesic expansion coefficient $\Theta$, bulk viscosity $\zeta$, shear viscosity $\eta$, momentum density $\pi^{A}$, shear tensor $\sigma^{AB}$ near the event horizon. If we identify (22) and (23) we get the following
@@
In particular, for the Schwarzschild geometry with $f=(1-\frac{2m}{r})$, we have:
@@
III.2 Doubly-modified Schwarzschild Metric
Johannsen-Psaltis black hole is a parametrically-deviated rotating black hole; before we work out its membrane construction, we would like to give the non-rotating version. This is because membrane paradigm analysis in the non-rotating limit ensures an easier detection of the transport coefficients, especially pressure according to truncation to the correct static limit. By taking the zero angular momentum limit to the JP black hole at hand we should be able to fix $\zeta=-\frac{1}{16\pi}$ condition [22]. This also implies that the non-rotating limit of the Johannsen-Psaltis black hole should have the same surface gravity as its doubly modified Schwarzschild counterpart. Now, we will present the doubly-modified Schwarzschild metric, we show that equality at the level of surface gravity brings another constraint on the deformation function and relates $\epsilon_{3}\simeq\alpha$. Firstly, let us introduce the metric and find its transport coefficients:
@@ III.2 Doubly-modified Schwarzschild Metric
where we have a Schwarzschild-like causal structure with two seemingly distinct additional hairs i.e. $\epsilon_{3}$ and $\alpha$ In the older version of the JP metric [9], one can see that the corrections to the Kerr metric were realized by applying the Newman-Janis algorithm to the modified Schwarzschild metric. Through this algorithm, the new hair is also complexified and naturally adapted to the Kerr-like metric. However, the metric in [10], the Johannsen metric, is already in Klein-Gordon separable form and the deformations only depend on radial coordinate $r$. This gives one the freedom to choose those new deformations in such a way that the metric has the same event horizon radius as the Kerr black hole. Hence, currently, its static counterpart has not been given in the literature. To be able to fix the bulk viscosity as $\zeta=-\frac{1}{16\pi}$, we choose its non-rotating limit to have the Schwarzschild radius.
One can realize that this metric has a similar causal structure as the Schwarzschild metric with the horizon at $r_{H}=2m$ as in the case of its rotating part has $r_{H}=m^{2}+\sqrt{m^{2}-a^{2}}$. Now, we can turn the paradigm’s machinery on and find its dual fluid correspondence. Firstly, one can directly choose its space-like normal vector as $n_{\mu}=\left\{0,\sqrt{\frac{h}{gF}},0,0\right\}$. Then the Parikh-Wilczek type decomposition reads as:
@@
and the 2D cross-section of the black hole metric is
@@
Note that in the limit of no-deformation, the stretched horizon stress tensor reduces to Schwarzschild’s membrane stress tensor given as
@@
By using the transport coefficients given in (24), we can proceed with the same analysis while choosing the nullness condition as:
@@
while the expansion, the shear tensor, and the surface gravity become:

### v2

@@ Abstract
The Johannsen-Psaltis (JP) black hole is a phenomenologically viable metric obtained by judiciously deforming the Kerr black hole such that the metric is asymptotically flat and causal and is consistent with the weak field, post-Newtonian tests of gravity; however, it has additional hairs besides its mass and angular momentum. As it deviates from the Kerr black hole in the strong coupling regime, it is a useful metric to test the Kerr hypothesis that states that all astrophysical black holes are of the Kerr type. Here we give a membrane description of this black hole that effectively amounts to replacing the observable part of the black hole with a fluid with thermal properties. A timelike membrane, a stretched horizon local in time is assumed to exist. This membrane is expected to approximate the null event horizon that is highly non-local in time. We derive the energy-momentum tensor of the fluid and all the transport coefficients using the action formulation to the membrane as advocated by Parikh and Wilczek. In the fluid description, one observes that the finiteness of the transport coefficients constrains the additional hairs of the Johannsen-Psaltis black hole. Analytically continuing the pressure of the fluid to all values of the radial coordinate $r$, one obtains interesting Van der Waals–type behavior of the pressure of the fluid which diverges at the radius of the outer ergosphere, lending support, from the membrane paradigm’s perspective to the claim that relativistic astrophysical jets are produced by the ergoregion of the black hole.
@@ I Introduction
Most of the deformations of the Kerr metric yield pathological results such as the existence of closed timelike curves, or naked singularities. Moreover, because of the powerful uniqueness theorems of the Kerr black hole [8] within general relativity, it is just in vain to modify the Kerr metric as a solution to Einstein’s equations: one must either consider alternative theories of gravity that allow viable deformations of the Kerr metric in the strong field regime or, perhaps better yet, without referring to any particular field equations, consider phenomenological metrics that judiciously deform metric. In the literature, one can find many works devoted to the first type of deformations, that is modified Kerr solutions (albeit almost always approximate solutions) to modified theories of Kerr gravity, and not so many phenomenological metrics except the ones constructed by Johannsen and Psaltis [10, 9]. Here we shall study these metrics in the context of the membrane paradigm approach to black holes which is an effective description of the black hole in terms of a fluid with various thermal properties. This is an exotic fluid (which may not be easily drinkable as it has a negative bulk viscosity) that mimics the true event horizon of the black hole. Our motivation to study the modified Kerr metric in the membrane paradigm was to understand if there are some constraints on the deformations parameters, and indeed there are as we shall see. In some sense, the theory is still a great “experimental tool” and a viable membrane paradigm that was quite useful in understanding various aspects of black holes in general relativity, helping us constrain the deformation parameter, indeed a new hair, in the parametrically deformed Kerr metric.
The layout of the paper is as follows: In Sec. II we give a detailed account of the membrane construction using the action formalism of Parikh-Wilczek [11]; Sec. III is the bulk of our paper: we construct the membranes for the JP metric after considering the membrane for its static limit. The computations presented here can be applied verbatim to any given stationary metric that forms an asymptotically flat deformation of the Kerr black hole.
@@ II.1 Basics of the membrane paradigm
The event horizon of a static or stationary black hole is a null hypersurface of codimension one with a degenerate metric. It is highly nonlocal in time and hence not observable to transient observers like us and is not a tangible concept for astrophysical purposes. To remedy this and to approximate the black hole as a local object in time, an effective description that makes use of timelike surfaces, not null surfaces is needed. For this purpose, a “stretched horizon” or “fluid bubble,” that is, a timelike surface [12] with electrical conductivity having a resistance of 377 $\Omega$, shear and bulk viscosities [13] arbitrarily close to the event horizon was introduced by Damour [13]. This pragmatic point of view, that is endowing a black hole with a timelike facade has proven to be very useful. A careful calculation on the horizon led to Ohm’s law [14], Joule’s law, and the nonrelativistic Navier-Stokes equation [15]. Early works of Damour, followed by Thorne and Price [16] showed a way to mimic the null horizon in a well-defined approximate way since the membrane is not null by nature, it has a nondegenerate Lorentzian metric which allows computations local in time. This approach to black hole dynamics was coined as the membrane paradigm of black holes [16]. More recently [11], a proper action formulation of the membrane was given that makes the computations rather straightforward, and hence we shall use this. See [17] for a recent review of these ideas.
@@ II.2 Quantities describing the fluid membrane
Let $\mathcal{H}$ be the event horizon of a stationary black hole, then there exists, by definition, a null geodesic vector field $\ell$ that generates $\mathcal{H}$. One can define a timelike stretched horizon $\mathcal{H}_{s}$ arbitrarily close to this event horizon. In the Arnowitt–Deser–Misner (ADM) decomposition of the metric [18], let $N$ be the lapse function that can be chosen such that in the limit $N\rightarrow 0$, the stretched horizon goes to the true horizon, that is $\mathcal{H}_{s}|_{N\rightarrow 0}=\mathcal{H}$ [15]. This limit in the geometry cannot be smooth and hence $N$ will be used as a regulator in various geometric objects as we shall see below.
Let $(\mathcal{M},g)$ be the $(3+1)$-dimensional total black hole spacetime and $(\mathcal{H}_{s},h)$ be a $(2+1)$-dimensional submanifold of $(\mathcal{M},g)$ with the pullback metric $h=\phi^{*}g$, which can also be considered as the projector $h^{\mu}\,_{\nu}:T_{p}(\mathcal{M})\longrightarrow T_{p}(\mathcal{H}_{s})$. A spacelike cross section of $\mathcal{H}_{s}$ is also a submanifold of dimension two, which we denote as $(\Sigma,\gamma)$. We can do differential geometry adapted to these two submanifolds, which boils down to a $(2+1+1)$ splitting of the full spacetime. Let $\nabla_{\mu}$ be the $g$-compatible covariant derivative and $D_{\mu}$ be the $h$-compatible covariant derivative while $\mathcal{D}_{\mu}$ be the $\gamma$-compatible covariant derivative [11]. Then, let $V^{\mu}\in T_{p}(\mathcal{M})$ and $n^{\mu}$ be a spacelike unit vector normal to $\mathcal{H}_{s}$, then defining the extrinsic curvature of $\mathcal{H}_{s}$ as $K_{\sigma\mu}:={h}{{}^{\sigma}}_{\mu}\nabla_{\sigma}n^{\mu}$, one has the identity
@@
Let $u^{\mu}$ be the unit normal timelike vector to $\mathcal{H}_{s}$ chosen to satisfy $u^{\mu}n_{\mu}=0$. $u^{\mu}$ can be considered to be the four velocity of a fiducial observer with proper time $\tau$. Here is the crucial part of the discussion: for the membrane paradigm to approximate the true horizon, a nullness constraint on the stretched horizon should be imposed; this means a change of character of $u^{\mu}$ and $n^{\mu}$ in the limits
@@
The relations we shall use for the membrane paradigm can be summarized as [11]
@@ II.3 An action formalism for the membrane paradigm
Parikh and Wilczek [11] gave an action formulation of the membrane that starts by modifying the usual variational principle
@@
where the surface term refers to the black hole boundary that now is represented effectively by the membrane. The first and second parts of the action variations are assumed to be zero individually [15].
@@ II.3.1 The gravitational membrane
where we assumed $G_{N}=1$, $c=1$ and the second term is the Gibbons-Hawking boundary term. Using the Palatini’s identity, one has [19]
@@
We choose the normal unit vector $n^{\mu}$ as outward pointing. Applying the Leibniz rule to the integrand gives
@@
Using the definition of the extrinsic curvature as given in the paragraph above (1), one finds the variation of the action as the Brown-York quasilocal stress tensor.
@@
Then, one has $t^{\text{stretched}}_{\mu\nu}=\frac{1}{8\pi}\left(Kh_{\mu\nu}-K_{\mu\nu}\right)\in T_{p}(\mathcal{H}_{s})\otimes T_{p}(\mathcal{H}_{s})$. On the stretched horizon $g^{\mu\nu}|_{\mathcal{H}_{s}}=h^{\mu\nu}$.To cancel the above nonzero boundary variation term, one must add the following:
@@
As it happens in electrodynamics where a surface charge induces a discontinuity in the field strength on the surface, $t^{\text{stretched}}_{\mu\nu}$ induces a discontinuity in the stretched horizon’s extrinsic curvature $K_{\mu\nu}$ [11]. This discontinuity creates a junction on the surface that can be identified as the Israel junction condition [20]:
@@
Hence the gravitational membrane acts like a fluid obeying Damour-Navier-Stokes equations on the spacelike cross section of the stretched horizon [21]. One can write ${K}{{}^{\mu}}_{\nu}$ in terms of the surface gravity $\kappa$ and extrinsic curvature ${k}{{}^{A}}_{B}$ of the spacelike section of $\mathcal{H}_{s}$. To this end, one has $\nabla_{\ell}\ell=\kappa_{r}\ell$ where $\kappa_{r}$ is the normalized surface gravity at the horizon, which is related to the surface gravity $\kappa$ as $\kappa_{r}=N\kappa$ [11].
@@
where $\sigma_{AB}$ is the shear tensor. Then finally, the stretched horizon stress tensor becomes [15]
@@ III.1 A membrane for the Schwarzschild-type geometry
where $f=1-\frac{2m(r)}{r}$. We can the metric as [^1] This construction follows [22].
@@
of which the trace is $K=\frac{1}{2}\frac{\partial_{r}f}{\sqrt{f}}+2\frac{\sqrt{f}}{r}$. In the $(2+1+1)$ splitting of the spacetime, the extrinsic curvature on $\mathcal{H}_{s}$ can be identified by choosing the lapse function $N=\sqrt{f}$ as a renormalization factor:
@@
where $k_{\mu\nu}=\gamma_{\mu A}\gamma_{\nu B}k^{AB}$ is the extrinsic curvature of the 2D surface and $\kappa$ is the surface gravity [22].
@@
Equations (19)–(21) can be combined to find the stress tensor in terms of given parameters:
@@
with energy density $\rho$, pressure $P$, null geodesic expansion coefficient $\Theta$, bulk viscosity $\zeta$, shear viscosity $\eta$, momentum density $\pi^{A}$, shear tensor $\sigma^{AB}$ near the event horizon. If we identify (22) and (23) we get the following:
@@
In particular, for the Schwarzschild geometry with $f=(1-\frac{2m}{r})$, we have
@@
III.2 Doubly modified Schwarzschild metric
Johannsen-Psaltis black hole is a parametrically deviated rotating black hole; before we work out its membrane construction, we would like to give the nonrotating version. This is because membrane paradigm analysis in the nonrotating limit ensures an easier detection of the transport coefficients, especially pressure according to truncation to the correct static limit. By taking the zero angular momentum limit to the JP black hole at hand we should be able to fix $\zeta=-\frac{1}{16\pi}$ condition [22]. This also implies that the nonrotating limit of the Johannsen-Psaltis black hole should have the same surface gravity as its doubly modified Schwarzschild counterpart. Now, we will present the doubly modified Schwarzschild metric, we show that equality at the level of surface gravity brings another constraint on the deformation function and relates $\epsilon_{3}\simeq\alpha$. First, let us introduce the metric and find its transport coefficients:
@@ III.2 Doubly modified Schwarzschild metric
where we have a Schwarzschild-like causal structure with two seemingly distinct additional hairs, i.e., $\epsilon_{3}$ and $\alpha$. In the older version of the JP metric [9], one can see that the corrections to the Kerr metric were realized by applying the Newman-Janis algorithm to the modified Schwarzschild metric. Through this algorithm, the new hair is also complexified and naturally adapted to the Kerr-like metric. However, the metric in [10], the Johannsen metric, is already in Klein-Gordon separable form and the deformations only depend on radial coordinate $r$. This gives one the freedom to choose those new deformations in such a way that the metric has the same event horizon radius as the Kerr black hole. Hence, currently, its static counterpart has not been given in the literature. To be able to fix the bulk viscosity as $\zeta=-\frac{1}{16\pi}$, we choose its nonrotating limit to have the Schwarzschild radius.
One can realize that this metric has a similar causal structure as the Schwarzschild metric with the horizon at $r_{H}=2m$ as in the case of its rotating part has $r_{H}=m^{2}+\sqrt{m^{2}-a^{2}}$. Now, we can turn the paradigm’s machinery on and find its dual fluid correspondence. Firstly, one can directly choose its spacelike normal vector as $n_{\mu}=\left\{0,\sqrt{\frac{h}{gF}},0,0\right\}$. Then the Parikh-Wilczek type decomposition reads as
@@
and the 2D cross section of the black hole metric is
@@
Note that in the limit of no deformation, the stretched horizon stress tensor reduces to Schwarzschild’s membrane stress tensor given as
@@
By using the transport coefficients given in (24), we can proceed with the same analysis while choosing the nullness condition as
@@
while the expansion, the shear tensor, and the surface gravity become